# Temporal Valued Constraint Satisfaction Problems

### v1

@@ 5.2. Expressibility of Valued Relations
Proof.
Note that $\operatorname{Cycl}(x,y,z)=\operatorname{Opt}(R_{\alpha,0,1}(x,y)+R_{\alpha,0,1}(y,z)+R_{\alpha,0,1}(z,x))$. Therefore, $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$. ∎
Lemma 5.15.
@@
Therefore, in both cases, $(<)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$. Since $(<)_{0}^{1}$ is equal to $R_{1,0,1}$, the statement follows from Lemma 5.14. ∎
Lemma 5.16.
@@
By Lemma 4.2, $\langle{\mathfrak{A}}\rangle$ contains $(<)_{0}^{\infty}$ or $(\neq)_{0}^{\infty}$. If $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$, then we are done. Assume therefore $(\neq)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$. Let $R$ be a valued relation of ${\mathfrak{A}}$ of arity $k$ such that there exists an orbit $O$ of the action of $\operatorname{Sym}(\mathbb{Q})$ on $\mathbb{Q}^{k}$ and $s,t\in O$ with $R(s)<R(t)$. Let $s\in O$ be such that $R(s)$ is minimal. Note that $O$ is not the orbit of constant tuples, because $\operatorname{Sym}(\mathbb{Q})$ is transitive.
@@
Consider the crisp relation $S\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ defined by
$S(x_{1},\dots,x_{k}):=\operatorname{Opt}\left(R(x_{1},\dots,x_{k})+\sum_{(p,q)\in E_{s}}(=)_{0}^{\infty}(x_{p},x_{q})+\sum_{(p,q)\in N_{s}}(\neq)_{0}^{\infty}(x_{p},x_{q})\right).$
@@
Clearly, $s\in S$. Note that a tuple $u\in\mathbb{Q}^{k}$ lies in $O$ if and only if $E_{s}\subseteq E_{u}$ and $N_{s}\subseteq N_{u}$. In particular, $S\subseteq O$. Since $R(s)<R(t)$, we have $t\notin S$. It follows that $S$ is not preserved by $\operatorname{Sym}(\mathbb{Q})$. Moreover, $S$ is not preserved by $\operatorname{const}$, because $O$ is not the orbit of constant tuples. By Theorem 5.2, the relational structure $(\mathbb{Q};S)$ admits a primitive positive definition of $\operatorname{Betw}$, $\operatorname{Cycl}$ or $\operatorname{Sep}$, or a primitive positive definition of $<$. Since $S\in\langle{\mathfrak{A}}\rangle$, the statement of the lemma follows. ∎
In Lemma 5.19 below we present a polynomial-time algorithm for VCSPs of valued structures ${\mathfrak{A}}$ improved by $\operatorname{lex}$ provided that ${\mathfrak{A}}$ cannot express any crisp relation that prevents tractability. In fact, to check whether the algorithm can be applied, it suffices to check whether a certain structure $\hat{{\mathfrak{A}}}$ with a finite signature has a tractable $\operatorname{CSP}$, instead of considering all relations in $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$. We define $\hat{{\mathfrak{A}}}$ below.
Definition 5.17.
Let $A$ be a set and let $R$ be a valued relation on $A$ of arity $k$. Let $\ell\in\mathbb{N}$, $\ell\leq k$, and let $\sigma\colon[k]\to[\ell]$ be a map. Then $R_{\sigma}$ is the valued relation on $A$ of arity $\ell$ defined by $R_{\sigma}(x_{1},\dots,x_{\ell})=R(x_{\sigma(1)},\dots,x_{\sigma(k)})$ for all $x_{1},\dots,x_{\ell}\in A$. If $S$ is a valued relation of some arity $\ell\leq k$ such that there exists $\sigma\colon[k]\to[\ell]$ and $S=R_{\sigma}$, we call $S$minor of $R$.
Let ${\mathfrak{A}}$ be a valued $\tau$-structure such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Then $\hat{{\mathfrak{A}}}$ denotes the relational structure with domain $\mathbb{Q}$ which contains the relations $\operatorname{Feas}(R^{\mathfrak{A}})$ and $\operatorname{Opt}((R^{{\mathfrak{A}}})_{\sigma})$ for every $R\in\tau$ of arity $k$, $\ell\leq k$, and $\sigma\colon[k]\to[\ell]$.
Note that $R_{\sigma}\in\langle(A;R)\rangle$ for every valued relation $R$ of arity $k$ and every $\sigma\colon[k]\to[\ell]$.
Remark 5.18.
Note that we do not need to include relations of the form $\operatorname{Feas}((R^{{\mathfrak{A}}})_{\sigma})$ in $\hat{{\mathfrak{A}}}$, because for every valued relation $R$ on $A$ of arity $k$ and $\sigma\colon[k]\to[\ell]$, we have
$\operatorname{Feas}(R_{\sigma})=\operatorname{Feas}(R)_{\sigma}$
and therefore $\operatorname{Feas}(R_{\sigma})\in\langle(A;\operatorname{Feas}(R))\rangle$.
Lemma 5.19.
Let ${\mathfrak{A}}$ be a valued structure over a finite signature such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Suppose that $\operatorname{lex}\in\operatorname{fPol}({\mathfrak{A}})$ and that $\hat{{\mathfrak{A}}}$ is preserved by one of the operations $\operatorname{const}$, $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$ or one of their duals. Then $\operatorname{VCSP}({\mathfrak{A}})$ is in P.
Proof.
If $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$, then $\operatorname{VCSP}({\mathfrak{A}})$ is in P by Lemma 3.1. We may therefore assume that $\operatorname{const}\notin\operatorname{fPol}({\mathfrak{A}})$. Let $R\in\langle{\mathfrak{A}}\rangle$ be of arity $k$. Since $\operatorname{lex}\in\operatorname{fPol}({\mathfrak{A}})$, it improves $R$. Therefore, for every injective tuple $s\in\mathbb{Q}^{k}$ and any $t\in\mathbb{Q}^{k}$, it holds that
$R(s)=R(\operatorname{lex}(s,t))\leq 1/2\cdot(R(s)+R(t)),$
Let $(\phi,u)$ be an instance of $\operatorname{VCSP}({\mathfrak{A}})$ with variable set $V=\{v_{1},\dots,v_{N}\}$. Note that $\phi$ interpreted over $\operatorname{Feas}({\mathfrak{A}})$ can be seen as an instance of $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$ where each summand $R(x_{1},\dots,x_{k})$ of $\phi$ is interpreted as $\operatorname{Feas}(R^{\mathfrak{A}})(x_{1},\dots,x_{k})$. By the assumption on $\hat{{\mathfrak{A}}}$, $\operatorname{Feas}({\mathfrak{A}})$ is preserved by one of the operations $\operatorname{const}$, $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals. Since $\operatorname{lex}\in\operatorname{fPol}({\mathfrak{A}})$, by Lemma 2.23 and Remark 2.22, $\operatorname{lex}\in\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))$. Since $\operatorname{const}\notin\operatorname{fPol}({\mathfrak{A}})$, by Lemma 4.2, $\operatorname{const}\notin\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))$. Then $\min$, $\operatorname{mx}$, $\operatorname{mi}$, or one of their duals preserves $\operatorname{Feas}({\mathfrak{A}})$, and, by Proposition 5.11, $\pi\pi\in\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))$ or $\pi\pi^{*}\in\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))$. Therefore, by Proposition 5.12, we always have that $\operatorname{ll}$ or $\operatorname{ll}^{*}$ preserves $\operatorname{Feas}({\mathfrak{A}})$. Hence, by Theorem 5.8, $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$ is solvable in polynomial time and we can use the polynomial-time algorithm from [9] based on the operation $\operatorname{ll}$ or $\operatorname{ll}^{*}$ to solve $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$. If $\phi$, viewed as a primitive positive formula, is not satisfiable over $\operatorname{Feas}({\mathfrak{A}})$, then the minimum of $\phi$ is above every rational threshold and $(\phi,u)$ is rejected. Otherwise, we may compute the set $E\subseteq V^{2}$ of all pairs $(x,y)$ such that $f(x)=f(y)$ in every solution of $f\colon V\to\mathbb{Q}$ of $\phi$ over $\operatorname{Feas}({\mathfrak{A}})$ (we may assume without loss of generality that $\operatorname{Feas}({\mathfrak{A}})$ contains the relation $(\neq)_{0}^{\infty}$; since $\operatorname{Feas}({\mathfrak{A}})$ is preserved by $\operatorname{lex}$ it suffices to test the unsatisfiability of $\phi\wedge x\neq y$ for each of these pairs). It follows from the definition of $\operatorname{Feas}$ that for every $g\colon V\to\mathbb{Q}$, if $\phi$ evaluates to a finite value in ${\mathfrak{A}}$ under the assignment $g$, then $g(x)=g(y)$ for every $(x,y)\in E$. Moreover, for every $(x,y)\in V^{2}\setminus E$, there exists $g\colon V\to\mathbb{Q}$ such that $\phi$ evaluates to a finite value under $g$ and $g(x)\neq g(y)$.
We create a new $\tau$-expression $\phi^{\prime}$ from $\phi$ by replacing each occurrence of $v_{j}$ by $v_{i}$ for every $(v_{i},v_{j})\in E$ such that $i<j$. Let $V^{\prime}$ be the set of variables of $\phi^{\prime}$. By the discussion above, the minimum for $\phi^{\prime}$ over ${\mathfrak{A}}$ equals the minimum for $\phi$. Moreover, for every $(x,y)\in(V^{\prime})^{2}$, there exists $g^{\prime}\colon V^{\prime}\to\mathbb{Q}$ such that $\phi^{\prime}$ evaluates to a finite value under $g^{\prime}$ and $g^{\prime}(x)\neq g^{\prime}(y)$. Let $\phi^{\prime}:=\phi^{\prime}_{1}+\dots+\phi^{\prime}_{n}$ where for every $j\in[n]$ the summand $\phi^{\prime}_{j}$ is an atomic $\tau$-expression. We execute the following procedure for each $j\in[n]$. Let $\phi^{\prime}_{j}=R(x_{1},\dots,x_{k})$. Let $y_{1}^{j},\dots,y_{\ell_{j}}^{j}$ be an enumeration of all distinct variables that appear in $\{x_{1},\dots,x_{k}\}$ and let $S_{j}$ be a valued relation of arity $\ell$ defined by $S_{j}(y_{1}^{j},\dots,y_{\ell_{j}}^{j})=R(x_{1},\dots,x_{k})$. Clearly, $S_{j}$ is a minor of $R$. Note that the relation $S_{j}$ might be different for every summand, even if they contain the same relation symbol $R$, due to possibly different variable identifications. Observe that, by the properties of $\phi^{\prime}$, there exists an injective tuple $s^{j}\in\mathbb{Q}^{\ell_{j}}$ such that $S_{j}(s^{j})$ is finite. Note that $S_{j}\in\langle{\mathfrak{A}}\rangle$, and let $m_{j}:=m_{S_{j}}$. By the discussion in the beginning of the proof, $S_{j}(s^{j})=m_{j}$ and $\operatorname{Opt}(S_{j})\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ consists of all tuples that evaluate to $m_{j}$ in $S_{j}$. Since $S_{j}$ attains only finitely many values, we can identify $m_{j}$ in polynomial time for every $j$.
Let ${\mathfrak{B}}$ be the relational structure with domain $\mathbb{Q}$ and relations $\operatorname{Opt}(S_{1}),\dots,\operatorname{Opt}(S_{n})$. Let $\psi$ be the instance of $\operatorname{CSP}({\mathfrak{B}})$ obtained from $\phi^{\prime}$ by replacing the summand $\phi^{\prime}_{j}$ by $\operatorname{Opt}(S_{j})(y_{1}^{j},\dots,y_{\ell_{j}}^{j})$ for all $j\in[n]$; all relations in $\psi$ are crisp and hence it can be seen as a primitive positive formula. Note that the variable set of $\psi$ is equal to $V^{\prime}$. By assumption, $\hat{{\mathfrak{A}}}$ is preserved by one of the operations $\operatorname{const}$, $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals and, in particular, ${\mathfrak{B}}$ is preserved by one of them. Hence, $\operatorname{CSP}({\mathfrak{B}})$ is in P by Theorem 5.8. Therefore, the satisfiability of $\psi$ over ${\mathfrak{B}}$ can be tested in polynomial time. We claim that if $\psi$ is unsatisfiable, then the minimum of $\phi$ is above every rational threshold and the algorithm rejects.
Finally, suppose that there exists a solution $h^{\prime}\colon V^{\prime}\to\mathbb{Q}$ to the instance $\psi$ of $\operatorname{CSP}({\mathfrak{B}})$. Then, for every $j\in[n]$, $\phi^{\prime}_{j}$ takes under $h^{\prime}$ the value $S_{j}(h^{\prime}(y_{1}^{j}),\dots,h^{\prime}(y_{\ell_{j}}^{j}))$. By the definition of $\operatorname{Opt}$, $(h^{\prime}(y_{1}^{j}),\dots,h^{\prime}(y_{\ell_{j}}^{j}))$ minimizes $S_{j}$ and therefore $h^{\prime}$ minimizes $\phi^{\prime}_{j}$. It follows that $h^{\prime}$ minimizes $\phi^{\prime}$ and that the cost of $\phi^{\prime}$ under $h^{\prime}$ is equal to $m_{1}+\dots+m_{n}$. Since the minimum of $\phi^{\prime}$ is equal to the minimum of $\phi$, the algorithm accepts if $m_{1}+\dots+m_{n}\leq u$ and rejects otherwise. This completes the algorithm and its correctness proof. It follows that $\operatorname{VCSP}({\mathfrak{A}})$ is in P. ∎
Lemma 5.20.
@@
Proof.
Let $R$ be a valued relation of ${\mathfrak{A}}$ of arity $k$ that attains at least two finite values. Let $m,\ell\in{\mathbb{Q}}$, with $m<\ell$, be the two smallest finite values attained by $R$. Let $t\in\mathbb{Q}^{k}$ be such that $R(t)=\ell$. Choose $s\in\operatorname{Opt}(R)$ so that $|(E_{s}\cap E_{t})\cup(O_{s}\cap O_{t})|$ is maximal (recall Definition 2.26). Clearly, $R(s)=m$.
Let $\sim\,\subseteq(\mathbb{Q}^{2})^{2}$ be the equivalence relation with the classes $=$, $<$, and $>$. Since $R(s)\neq R(t)$, there exist distinct $i,j$ such that $(s_{i},s_{j})\not\sim(t_{i},t_{j})$. For the sake of notation, assume that $(i,j)=(1,2)$. Let
$\phi(x_{1},x_{2}):=\min_{x_{3},\dots,x_{k}}\left(R(x_{1},\dots,x_{k})+\sum_{(p,q)\in E_{s}\cap E_{t}}(=)_{0}^{\infty}(x_{p},x_{q})+\sum_{(p,q)\in O_{s}\cap O_{t}}(<)_{0}^{\infty}(x_{p},x_{q})\right).$
@@
By the construction, $S\in\langle{\mathfrak{A}}\rangle$, $S(x,y)=0$ for $(x,y)\sim(s_{1},s_{2})$, and $S(x,y)=1$ for $(x,y)\sim(t_{1},t_{2})$. Note that $\operatorname{Aut}(\mathbb{Q};<)$ has three orbits of pairs, two of which are represented by $(s_{1},s_{2})$ and $(t_{1},t_{2})$. Let $(u_{1},u_{2})\in\mathbb{Q}^{2}$ be a representative of the third orbit and let $\alpha=S(u_{1},u_{2})$. It follows that $S$ is equal to one of the relations $R_{0,1,\alpha}$, $R_{0,\alpha,1}$, $R_{1,0,\alpha}$, $R_{1,\alpha,0}$, $R_{\alpha,0,1}$ or $R_{\alpha,1,0}$. By the choice of $m$ and $\ell$, we have that $\alpha=0$ or $\alpha\geq 1$. By Lemma 5.15, this implies that $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$ unless $S=R_{1,0,0}$, $S=R_{1,0,\infty}$, or $S=R_{1,\infty,0}$. Since $R_{1,0,0}$ is equal to $(\neq)_{0}^{1}$ and $R_{1,0,\infty}(x,y)=R_{1,\infty,0}(y,x)$, the statement follows. ∎
Lemma 5.21.
@@
Let ${\mathfrak{A}}$ be a valued structure with $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};<)$ and $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$. Suppose that $\operatorname{lex}\notin\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ and that ${\mathfrak{A}}$ is not essentially crisp. Then $\langle{\mathfrak{A}}\rangle$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, or $\operatorname{Dis}$.
Proof.
@@
Let ${\mathfrak{A}}^{\prime}:=(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. Note that $\operatorname{const}\notin\operatorname{Pol}({\mathfrak{A}}^{\prime})$, because $\operatorname{const}$ does not preserve $(<)_{0}^{\infty}$. If $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, or $\operatorname{Dis}$, then we are done. Suppose that this is not the case. Then $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals by Theorem 5.8. Suppose first that $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, or $\operatorname{ll}$. Since $\operatorname{lex}\notin\operatorname{Pol}({\mathfrak{A}}^{\prime})$, we have $\operatorname{ll}\notin\operatorname{Pol}({\mathfrak{A}}^{\prime})$ (Lemma 5.6). By Proposition 5.11, $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\pi\pi$. By Theorem 5.13, ${\mathfrak{A}}^{\prime}$ primitively positively defines, equivalently, contains the relation $R^{\operatorname{mix}}$. By Lemma 5.20, we have that $\langle{\mathfrak{A}}\rangle$ contains $\operatorname{Cycl}$, $(\neq)_{0}^{1}$, or $R_{1,0,\infty}$. If $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$, then we are done. Suppose therefore that $(\neq)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$ or $R_{1,0,\infty}\in\langle{\mathfrak{A}}\rangle$. Note that for every $x,y\in\mathbb{Q}$, we have
$(<)_{0}^{1}(x,y)$ $=\min_{z}\big{(}R^{\operatorname{mix}}(y,z,x)+(\neq)_{0}^{1}(y,z)\big{)}=\min_{z}\big{(}R^{\operatorname{mix}}(y,z,x)+R_{1,0,\infty}(y,z)\big{)}.$

### v2

@@ 5.3 Expressibility of hard valued relations
Proof 11
Note that
$\operatorname{Cycl}(x,y,z)=\operatorname{Opt}(R_{\alpha,0,1}(x,y)+R_{\alpha,0,1}(y,z)+R_{\alpha,0,1}(z,x)).$
Therefore, $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$. $\qed$
Lemma 49
@@
Therefore, in both cases, $(<)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$. Since $(<)_{0}^{1}$ is equal to $R_{1,0,1}$, the statement follows from Lemma 48. $\qed$
The following lemma shows that if $\mathfrak{A}$} is not an equality valued structure and has a tractable VCSP, then $\operatorname{const}\notin\operatorname{fPol}({\mathfrak{A}})$ implies that $(<)_{0}^{\infty}$ is pp-expressible in $\mathfrak{A}$}. The proof is based on isolating an expressible relation not preserved by $\operatorname{Sym}(\mathbb{Q})$ and Theorem 35.
Lemma 50
@@
By Lemma 28, $\langle{\mathfrak{A}}\rangle$ contains $(<)_{0}^{\infty}$ or $(\neq)_{0}^{\infty}$. If $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$, then we are done. Assume therefore $(\neq)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$. Let $R$ be a valued relation of $\mathfrak{A}$} of arity $k$ such that there exists an orbit $O$ of the action of $\operatorname{Sym}(\mathbb{Q})$ on $\mathbb{Q}^{k}$ and $s,t\in O$ with $R(s)<R(t)$. Let $s\in O$ be such that $R(s)$ is minimal. Note that $O$ is not the orbit of constant tuples, because $\operatorname{Aut}(\mathbb{Q};<)$ is transitive.
@@
Consider the crisp relation $S\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ defined by the following expression with the free variables $x_{1},\dots,x_{k}$.
$\operatorname{Opt}\left(R(x_{1},\dots,x_{k})+\sum_{(p,q)\in E_{s}}(=)_{0}^{\infty}(x_{p},x_{q})+\sum_{(p,q)\in N_{s}}(\neq)_{0}^{\infty}(x_{p},x_{q})\right)$
@@
Clearly, $s\in S$. Note that a tuple $u\in\mathbb{Q}^{k}$ lies in $O$ if and only if $E_{s}\subseteq E_{u}$ and $N_{s}\subseteq N_{u}$. In particular, $S\subseteq O$. Since $R(s)<R(t)$, we have $t\notin S$. It follows that $S$ is not preserved by $\operatorname{Sym}(\mathbb{Q})$. Moreover, $S$ is not preserved by $\operatorname{const}$, because $O$ is not the orbit of constant tuples. By Theorem 35, the relational structure $(\mathbb{Q};S)$ admits a primitive positive definition of $\operatorname{Betw}$, $\operatorname{Cycl}$ or $\operatorname{Sep}$, or a primitive positive definition of <<. Since $S\in\langle{\mathfrak{A}}\rangle$, the statement of the lemma follows. $\qed$
As a next step, we use the assumption that $\mathfrak{A}$} is not essentially crisp and the relation $(<)_{0}^{\infty}$ to express some useful valued relations.
Lemma 51
@@
Proof 14
Let $R$ be a valued relation of $\mathfrak{A}$} of arity $k$ that attains at least two finite values. Let $m,\ell\in{\mathbb{Q}}$, with $m<\ell$, be the two smallest finite values attained by $R$. Let $t\in\mathbb{Q}^{k}$ be such that $R(t)=\ell$. Choose $s\in\operatorname{Opt}(R)$ so that $|(E_{s}\cap E_{t})\cup(O_{s}\cap O_{t})|$ is maximal (recall definition 22). Clearly, $R(s)=m$.
Let $\sim\,\subseteq(\mathbb{Q}^{2})^{2}$ be the equivalence relation with the classes ==, <<, and >>. Since $R(s)\neq R(t)$, there exist distinct $i,j$ such that $(s_{i},s_{j})\not\sim(t_{i},t_{j})$. For the sake of notation, assume that $(i,j)=(1,2)$. Let $\phi$ be the following expression with the free variables $x_{1},x_{2}$.
$\min_{x_{3},\dots,x_{k}}\left(R(x_{1},\dots,x_{k})+\sum_{(p,q)\in E_{s}\cap E_{t}}(=)_{0}^{\infty}(x_{p},x_{q})+\sum_{(p,q)\in O_{s}\cap O_{t}}(<)_{0}^{\infty}(x_{p},x_{q})\right)$
@@
By the construction, $S\in\langle{\mathfrak{A}}\rangle$, $S(x,y)=0$ for $(x,y)\sim(s_{1},s_{2})$, and $S(x,y)=1$ for $(x,y)\sim(t_{1},t_{2})$. Note that $\operatorname{Aut}(\mathbb{Q};<)$ has three orbits of pairs, two of which are represented by $(s_{1},s_{2})$ and $(t_{1},t_{2})$. Let $(u_{1},u_{2})\in\mathbb{Q}^{2}$ be a representative of the third orbit and let $\alpha=S(u_{1},u_{2})$. It follows that $S$ is equal to one of the relations $R_{0,1,\alpha}$, $R_{0,\alpha,1}$, $R_{1,0,\alpha}$, $R_{1,\alpha,0}$, $R_{\alpha,0,1}$ or $R_{\alpha,1,0}$. By the choice of $m$ and $\ell$, we have that $\alpha=0$ or $\alpha\geq 1$. By Lemma 49, this implies that $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$ unless $S=R_{1,0,0}$, $S=R_{1,0,\infty}$, or $S=R_{1,\infty,0}$. Since $R_{1,0,0}$ is equal to $(\neq)_{0}^{1}$ and $R_{1,0,\infty}(x,y)=R_{1,\infty,0}(y,x)$, the statement follows. $\qed$
The following lemma relates expressibility of hard relations to not having $\operatorname{ll}$ and $\operatorname{ll}^{*}$ in $\operatorname{Pol}(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. The essential ingredient of the proof is Theorem 42.
Lemma 52
@@
Let $\mathfrak{A}$} be a valued structure such that $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};<)$ and $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$. Suppose that $\operatorname{ll},\operatorname{ll}^{*}\notin\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ and that $\mathfrak{A}$} is not essentially crisp. Then $\langle{\mathfrak{A}}\rangle$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, or $\operatorname{Dis}$.
Proof 15
@@
Let ${\mathfrak{A}}^{\prime}:=(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. Note that $\operatorname{const}\notin\operatorname{Pol}({\mathfrak{A}}^{\prime})$, because $\operatorname{const}$ does not preserve $(<)_{0}^{\infty}$. If $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, or $\operatorname{Dis}$, then we are done. Suppose that this is not the case. Then $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, or one of their duals by Theorem 44 and the assumptions of the lemma. Suppose first that $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$. By Proposition 41, $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\pi\pi$. By Theorem 42, $\mathfrak{A}$$^{\prime}$ primitively positively defines, equivalently, contains the relation $R^{\operatorname{mix}}$. By Lemma 51, we have that $\langle{\mathfrak{A}}\rangle$ contains $\operatorname{Cycl}$, $(\neq)_{0}^{1}$, or $R_{1,0,\infty}$. If $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$, then we are done. Suppose therefore that $(\neq)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$ or $R_{1,0,\infty}\in\langle{\mathfrak{A}}\rangle$. Note that for every $x,y\in\mathbb{Q}$, we have
$(<)_{0}^{1}(x,y)$ $=\min_{z}\big(R^{\operatorname{mix}}(y,z,x)+(\neq)_{0}^{1}(y,z)\big)$
$=\min_{z}\big(R^{\operatorname{mix}}(y,z,x)+R_{1,0,\infty}(y,z)\big).$